2. The information given below has been taken from the costing records of an Engineering works in respect of Job No. 303.
Materials Sh. 41,100
Wages:
Department A – 60 Hours at Sh.300 per hour
Department B – 40 Hours at Sh.200 per hour
Department C – 20 Hours at Sh.50 per hour
Overhead expenses for these three departments were estimated as follows:
Variable overheads:
Department A – Sh.50,000 for 5,000 labour hours
Department B – Sh.30,000 for 1,500 labour hours
Department C – Sh.20,000 for 500 labour hours
Fixed overheads:
Estimated at Sh.200,000 for 10,000 normal working hours.
Required:
i) Calculate the cost of job 303
1
Expert's answer
2020-12-10T14:28:43-0500

SOLUTION TO QUESTION Q149858

COST SHEET JOB NUMBER NO 303

Direct materials Sh. 41,100

Add wages: Department

A =60 hrs*Sh. 300/hr= Sh. 18,000

B =40hrs*Sh. 200/hr= Sh. 8,000

C=20hrs*Sh.50/hr= Sh. 1,000

Total wages Sh. 27,000

Prime cost(Direct material+Wages) Sh. 68,100

Add Overhead variables: Department

A =60 hrs*Sh. 10/hr= Sh. 600

B =40hrs*Sh. 20/hr= Sh. 800

C=20hrs*Sh.40/hr= Sh. 800

Total overhead costs Sh. 2,200

Fixed costs overhead (60+40+20)*Sh. 20/hr Sh. 2,400

Total job cost Sh. 72,700